# Cost-Aware Contrastive Routing for LLMs

### v1

@@ 4.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines. Specifically, we include UMR Jitkrittum et al. (2025), a recent technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling Li (2025); Agrawal and Goyal (2012), which frames routing as a bandit exploration-exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing Hu et al. (2024), a strategy that selects models by explicitly considering the cost-accuracy Pareto frontier; and two extreme baselines: Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle Jitkrittum et al. (2025), which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric-softmax gating methods inspired by mixture-of-experts architectures Ong et al. (2024) and SoftMoE Puigcerver et al. (2024), which models router decisions via differentiable soft gating functions.
@@
We train our router and evaluate it on three datasets: EmbedLLM Zhuang et al. (2024), MixInstruct Jiang et al. (2023), and RouterBench Hu et al. (2024). For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)). For RouterBench, we sample 192 probes from its training set. We compute perplexity-based descriptors on RouterBench and use GPT-2 Radford et al. (2019). On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric. For MixInstruct, we employ exponentiated BARTScore Yuan et al. (2021) as the evaluation metric, following the approach in prior work Jitkrittum et al. (2025); Jiang et al. (2023).
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2Reimers and Gurevych (2019) model as the embedding backbone across all experiments. Our trainable router component is a two-layer MLP which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset. For the cost spectrum loss (Equation (8)), we set the number of cost bands $K=5$ and the negative cost penalty $\gamma=0.2$. The hyperparameters for band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. See D.1 for full details.
@@
We evaluate each routing strategy using a deferral curve Jitkrittum et al. (2025) which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following Jitkrittum et al. (2025) we employ evaluation metrics including Area Under the Deferral Curve (AUDC), peak accuracy and Query-Normalized Cost (QNC), the minimum relative cost required to match the performance of the most accurate tested LLM.
@@ 4.2 Results
EmbedLLM Mix-Instruct RouterBench
@@
UMR Jitkrittum et al. (2025) 0.515 58.61 0.562 0.049 10.35 0.050 0.568 0.487 0.615
Thompson Agrawal and Goyal (2012) 0.472 48.81 0.514 0.044 10.09 0.045 0.622 1.634 0.787
Soft-MoE Puigcerver et al. (2024) 0.404 70.00 0.577 0.030 10.09 0.045 0.599 1.659 0.794
Parametric Ong et al. (2024) 0.506 70.00 0.555 0.039 12.00 0.052 0.691 1.658 0.794
Pareto-Random Hu et al. (2024) 0.369 16.03 0.393 0.036 6.16 0.043 0.5172 0.589 0.545
@@
Table 1 presents deferral-curve metrics across the benchmarks. Our CSCR consistently outperforms all relevant baselines, achieving the highest AUDC and demonstrating competitive or superior peak accuracy. Notably, it attains lower QNC, indicating more cost-effective routing decisions. These results underscore the effectiveness of our cost-aware router learning approach in balancing performance and inference cost. The Oracle router, which retrospectively selects the optimal LLM for each query, establishes an upper bound for performance. The benchmarks are dominated by lower cost experts, hence the lower QNC for random baselines.
@@ 4.2.1 Generalization to New LLMs
Soft-MoE 0.4109 18.282 0.485
Pareto-Random 0.3812 15.662 0.407
@@
Table 2: Deferral-curve metrics on new LLMs. CSCR shows better robustness to new LLMs.
@@
Soft-MoE 0.4207 70.017 0.555
Pareto-Random 0.3676 17.417 0.396
@@
Table 3: Deferral-curve metrics on OOD prompts. CSCR shows superior robustness.
See Appendix D.6 for results on statistical significance.
We also evaluate the robustness of our method and baselines in scenarios where new LLMs are introduced during testing. Specifically, we select two-thirds of the EmbedLLM training models to train our router and the baselines following Jitkrittum et al. (2025), using only responses from these selected models. Table 2 summarizes the performance metrics, while Figure 3 illustrates the corresponding deferral curves. Testing is conducted exclusively on the unseen LLM pool. Our results indicate that our approach exhibits superior robustness under these conditions.
@@ 4.3.2 Cost-Aware Training
Replacing the vanilla InfoNCE loss with our cost-spectrum variant significantly increases the AUDC (0.342 $\rightarrow$ 0.495) and raises the peak attainable accuracy from 36% to 54% as shown in Table 5. The trade-off is a higher Quality-Neutral Cost QNC, meaning the router now leans more on expensive but accurate models; however, the large AUDC gain shows that, for any realistic cost budget, users receive better accuracy-per-dollar overall. This confirms our discussion in Section 3.4, that explicitly teaching the encoder to respect the cost hierarchy of experts is crucial.
@@
Table 5: Effect of cost-aware training: injecting cost awareness into the contrastive loss prevents the router from concentrating on cheap experts, and boosts the AUDC and peak accuracy of the router. The trade-off is a higher QNC.
@@ 4.3.3 Cost-Spectrum Granularity
Figure 6: Deferral-curves across different numbers of cost bands.
Figure 6 shows that performance varies with the number of cost bands. Using 5 bands yields the best results, with the highest AUDC and Peak accuracy, suggesting a good balance between flexibility and generalization. Fewer bands (e.g., 2) limit routing precision, while too many bands (e.g., 15 or 20) degrade performance, likely due to over-fragmentation and increased decision noise. This highlights the importance of tuning the number of bands per dataset to avoid both under- and overfitting.
See Appendix D.2 for more ablations and detailed discussion.
@@ 5 Theoretical Analysis
Set-up.
Let $\mathcal{X}$ be the space of prompts and $\mathcal{H}=\{h_{1},\dots,h_{M}\}$ a fixed pool of experts with per-call cost $c(h)\in\mathbb{R}_{>0}$. For a query $x\in\mathcal{X}$ and ground-truth $y$ we write $\ell(x,y,h)=\mathbbm{1}\![h(x)\neq y]$ for the 011 loss and $\gamma(x,h)=\mathbb{P}_{y\mid x}[\ell(x,y,h)=1]$ for the Bayes error. The cost–adjusted risk of a router $r$ is
@@
where $\lambda\!\in\!\mathbb{R}_{\geq 0}$ trades accuracy for cost Elkan (2001).
Our router embeds queries via $\Phi_{q}:\mathcal{X}\!\to\!\mathbb{R}^{d}$ and experts via $E=\bigl{[}e_{1};\dots;e_{M}\bigr{]}\!\in\!\mathbb{R}^{M\times d}$, using either (i) logits fingerprints (EmbedLLM, Mix-Instruct) or (ii) perplexity fingerprints (RouterBench). Given a query $x$, the $k$-NN rule selects
@@
5.1 Excess-risk of cost–spectrum k-NN
@@
where $C$ depends only on $L$ and the diameter of $\Phi_{q}(\mathcal{X})$, and $r^{\star}$ is the Bayes-optimal rule $r^{\star}(x)=\mathop{\rm arg\,min}_{m}\bigl{[}\gamma(x,h_{m})+\lambda c(h_{m})\bigr{]}$.
The proof follows the classical $k$-NN bound of Luxburg and Bousquet (2004); Biau and Devroye (2015) with an extra $\lambda c(h)$ term that is constant w.r.t. $x$ and therefore preserves the rate.
@@ 5.2 Consistency of Cost–Spectrum InfoNCE
Fingerprints live on $\mathbb{S}^{d-1}$, so dot products are scaled similarities $S_{im}=\tfrac{q_{i}^{\top}e_{m}}{\tau_{k}}$ with band-dependent temperature $\tau_{k}$ (Equation (7)). Let $\beta_{0}\!=\!0<\!\beta_{1}<\!\dots<\!\beta_{K}\!=\!1$ partition costs into bands $\mathcal{B}_{k}=\{m:\;c(h_{m})\!\in\![\beta_{k},\beta_{k+1})\}$, and denote $\mathcal{P}_{ik}=\mathcal{P}(i)\cap\mathcal{B}_{k}$ the correct experts for query $i$ that fall in band $k$. The Cost-Spectrum InfoNCE loss for a single query $i$ is
@@ 5.3 Discussion
Theorem 5.2 guarantees that if query–expert descriptors are Lipschitz, CSCR converges to the Bayes-optimal router at the usual $k$-NN rate. Lemma 5.3 justifies the specific form of our InfoNCE objective. Sec. 4 verifies these claims on three benchmarks. See proofs in Appendix B.
@@ A.1 Enhancing and Optimizing LLMs
Large language models (LLMs) have demonstrated remarkable capabilities across diverse NLP tasks Radford et al. (2019); Brown et al. (2020). To further improve their performance and efficiency, numerous strategies have been proposed.
@@
Enhancement approaches targeting individual LLMs include fine-tuning Rafailov et al. (2023), prompting strategies like Chain-of-Thought (CoT) Zhou et al. (2023); Wang et al. (2022), Tree-of-Thoughts Yao et al. (2023), and inference-acceleration techniques such as early exiting Teerapittayanon et al. (2016); Zhou et al. (2020); Schuster et al. (2022) and speculative decoding Stern et al. (2018); Sun et al. (2024); Chen et al. (2023a); Leviathan et al. (2023); Cai et al. (2024b). Additionally, Mixture-of-Experts (MoE) architectures Jacobs et al. (1991); Jordan and Jacobs (1993); Shazeer et al. (2017); Fedus et al. (2022); Zhou et al. (2022); Jiang et al. (2024) route inputs through sparse sub-models or "experts," reducing cost while retaining performance. However, these methods typically operate within a single LLM’s structure and may not generalize to multi-model scenarios.
@@
Fusion strategies synthesize outputs from multiple LLMs to improve output quality Ravaut et al. (2022); Jiang et al. (2023); Guha et al. (2024); Wang et al. (2024); Lu et al. (2024b). Fusion approaches often rely on unsupervised metrics Zhang* et al. (2020); Sellam et al. (2020); Yuan et al. (2021) or ensemble voting to determine the final output Lee et al. (2023b). A related but distinct technique is model merging Lu et al. (2024a), where weights from multiple pre-trained or fine-tuned models are combined—either directly via methods like weight averaging Wortsman et al. (2022), Task Arithmetic Ilharco et al. (2023), or Fisher merging Matena and Raffel (2022)—or using more sophisticated techniques like TIES Yadav et al. (2023), AdaMerging Yang et al. (2023), and ZipIt Stoica et al. (2024).
@@
In contrast, cascading invokes models sequentially—often ordered by computational cost—and halts once a satisfactory output is generated Chen et al. (2023b); Yue et al. (2023); Chen et al. (2023b); Gupta et al. (2024). Such approaches strike a balance between quality and efficiency, making them particularly attractive in production settings.
@@ A.2 LLM Routing
Non-predictive methods generate outputs from one or more models before making a selection. FrugalGPT Chen et al. (2023b) exemplifies this category, using a sequential strategy and a response quality threshold to minimize cost. Other works adopt layered inference architectures to escalate hard queries to more powerful models Wang et al. (2023), or leverage cascades with self-verification Madaan et al. (2023); Yue et al. (2023); Lee et al. (2023a); Lu et al. (2023).
@@
In contrast, predictive routing aims to select the best model *before* any inference is performed. Strategies include supervised learning Shnitzer et al. (2023), reward-model-based routing Hari and Thomson (2023); Lu et al. (2023), and meta-models trained to predict LLM performance given an input Šakota et al. (2023). Router models vary widely in implementation, including neural networks Ding et al. (2024a); Šakota et al. (2024); Chen et al. (2024); Aggarwal et al. (2024), $k$-nearest neighbors Hu et al. (2024); Shnitzer et al. (2023); Stripelis et al. (2024); Lee et al. (2023a), matrix factorization Ong et al. (2024); Zhuang et al. (2024); Li (2025), and graph neural networks Feng et al. (2024). Others incorporate model-specific tokens or train across multiple domains Devvrit et al. (2024); Cai et al. (2024a).
@@
Routing and cascading are grounded in broader literature, including selective classification Geifman and El-Yaniv (2019); Narasimhan et al. (2024), learning to defer Madras et al. (2018), and learning to reject Chow (1970); Bartlett and Wegkamp (2008); Cortes et al. (2016). Several works explore supervision levels Lu et al. (2023); Zhao et al. (2024b), robustness Dann et al. (2024); Montreuil et al. (2025); Shafran et al. (2025), and evaluation frameworks for routers Hendy et al. (2023); Hari and Thomson (2023).

### v2

@@ 4.1 Experimental Settings
We compare our proposed method against a comprehensive set of baselines. Specifically, we include UMR [40], a recent technique that clusters prompt embeddings to route queries to LLM pools efficiently; Thompson Sampling [48, 2], which frames routing as a bandit exploration-exploitation problem to balance cost and accuracy dynamically; Pareto-optimal routing [34], a strategy that selects models by explicitly considering the cost-accuracy Pareto frontier; and two extreme baselines: Random, which selects models uniformly at random to represent naive routing without intelligent selection, and Oracle [40], which always selects the most accurate model at the lowest possible cost and thus represents a theoretical performance ceiling. Additionally, we evaluate against parametric-softmax gating methods inspired by mixture-of-experts architectures [62] and SoftMoE [64], which models router decisions via differentiable soft gating functions.
@@
We train our router and evaluate it on three datasets: EmbedLLM [105], MixInstruct [39], and RouterBench [34]. For EmbedLLM and MixInstruct, we sample 192 probes from their respective validation sets. Each probe is processed to extract logit-based descriptors by capturing the top $K=256$ tokens over a horizon of $T=10$ tokens (Equation (2)). For RouterBench, we sample 192 probes from its training set. We compute perplexity-based descriptors on RouterBench and use GPT-2 [66]. On both EmbedLLM and RouterBench, we use binary accuracy as the per-sample evaluation metric. For MixInstruct, we employ exponentiated BARTScore [97] as the evaluation metric, following the approach in prior work [40, 39].
@@
We use a frozen sentence-transformers/all-MiniLM-L6-v2[70] model as the embedding backbone across all experiments. Our trainable router component is a two-layer MLP which projects prompt embeddings into the expert descriptor space. We train our contrastive router on the training splits of each dataset. For the cost spectrum loss (Equation (8)), we set the number of cost bands $K=5$ and the negative cost penalty $\gamma=0.2$. The hyperparameters for band-specific temperatures (Equation (7)) are set as $\alpha=0.25$ and $\tau_{\min}=0.05$. See D.1 for full details.
@@
We evaluate each routing strategy using a deferral curve [40] which plots the average response quality against the total inference cost. Sweeping the routing penalty parameter $\lambda$ over the interval $\lambda\!\in\![0,\lambda_{\max}]$ (Equation (9)) traces the deferral curve. For the EmbedLLM and MixInstruct datasets, we define the cost of processing a prompt as the number of parameters in the LLM, a proxy for computational resources and latency. In the case of RouterBench, we utilize the actual API call costs in USD, as provided in the dataset. Following [40] we employ evaluation metrics including Area Under the Deferral Curve (AUDC), peak accuracy and Query-Normalized Cost (QNC), the minimum relative cost required to match the performance of the most accurate tested LLM.
@@ 4.2 Results
EmbedLLM Mix‑Instruct RouterBench
@@
UMR [40] 0.515 58.61 0.562 0.049 10.35 0.050 0.568 0.487 0.615
Thompson [2] 0.472 48.81 0.514 0.044 10.09 0.045 0.622 1.634 0.787
Soft‑MoE [64] 0.404 70.00 0.577 0.030 10.09 0.045 0.599 1.659 0.794
Parametric [62] 0.506 70.00 0.555 0.039 12.00 0.052 0.691 1.658 0.794
Pareto‑Random [34] 0.369 16.03 0.393 0.036 6.16 0.043 0.5172 0.589 0.545
@@
Table 1 presents deferral-curve metrics across the benchmarks. Our CSCR consistently outperforms all relevant baselines, achieving the highest AUDC and demonstrating competitive or superior peak accuracy. Notably, it attains lower QNC, indicating more cost-effective routing decisions. These results underscore the effectiveness of our cost-aware router learning approach in balancing performance and inference cost. The Oracle router, which retrospectively selects the optimal LLM for each query, establishes an upper bound for performance. The benchmarks are dominated by lower cost experts, hence the lower QNC for random baselines. See Appendix D.6 for results on statistical significance.
@@ 4.2.1 Generalization to New LLMs
Soft‑MoE 0.4109 18.282 0.485
Pareto‑Random 0.3812 15.662 0.407
@@
Table 2: Deferral‑curve metrics on new LLMs. CSCR shows better robustness to new LLMs.
@@
Soft‑MoE 0.4207 70.017 0.555
Pareto‑Random 0.3676 17.417 0.396
@@
Table 3: Deferral‑curve metrics on OOD prompts. CSCR shows superior robustness.
We also evaluate the robustness of our method and baselines in scenarios where new LLMs are introduced during testing. Specifically, we select two-thirds of the EmbedLLM training models to train our router and the baselines following [40], using only responses from these selected models. Table 2 summarizes the performance metrics, while Figure 3 illustrates the corresponding deferral curves. Testing is conducted exclusively on the unseen LLM pool. Our results indicate that our approach exhibits superior robustness under these conditions.
@@ 4.3.2 Cost-Aware Training
Replacing the vanilla InfoNCE loss with our cost‑spectrum variant significantly increases the AUDC (0.342 $\rightarrow$ 0.495) and raises the peak attainable accuracy from 36% to 54% as shown in Table 5. The trade‑off is a higher Quality‑Neutral Cost QNC, meaning the router now leans more on expensive but accurate models; however, the large AUDC gain shows that, for any realistic cost budget, users receive better accuracy‑per‑dollar overall. This confirms our discussion in Section 3.4, that explicitly teaching the encoder to respect the cost hierarchy of experts is crucial.
@@
Table 5: Effect of cost‑aware training: injecting cost awareness into the contrastive loss prevents the router from concentrating on cheap experts, and boosts the AUDC and peak accuracy of the router. The trade‑off is a higher QNC.
@@ 4.3.3 Cost-Spectrum Granularity
Figure 6: Deferral‑curves across different numbers of cost bands.
Figure 6 shows that performance varies with the number of cost bands. Using 5 bands yields the best results, with the highest AUDC and Peak accuracy, suggesting a good balance between flexibility and generalization. Fewer bands (e.g., 2) limit routing precision, while too many bands (e.g., 15 or 20) degrade performance, likely due to over-fragmentation and increased decision noise. This highlights the importance of tuning the number of bands per dataset to avoid both under- and overfitting. See Appendix D.2 for more ablations and detailed discussion.
@@ 5 Theoretical Analysis
Set‑up.
Let $\mathcal{X}$ be the space of prompts and $\mathcal{H}=\{h_{1},\dots,h_{M}\}$ a fixed pool of experts with per‑call cost $c(h)\in\mathbb{R}_{>0}$. For a query $x\in\mathcal{X}$ and ground‑truth $y$ we write $\ell(x,y,h)=\mathbbm{1}\![h(x)\neq y]$ for the 011 loss and $\gamma(x,h)=\mathbb{P}_{y\mid x}[\ell(x,y,h)=1]$ for the Bayes error. The cost–adjusted risk of a router $r$ is
@@
where $\lambda\!\in\!\mathbb{R}_{\geq 0}$ trades accuracy for cost [23].
Our router embeds queries via $\Phi_{q}:\mathcal{X}\!\to\!\mathbb{R}^{d}$ and experts via $E=\bigl[e_{1};\dots;e_{M}\bigr]\!\in\!\mathbb{R}^{M\times d}$, using either (i) logits fingerprints (EmbedLLM, Mix‑Instruct) or (ii) perplexity fingerprints (RouterBench). Given a query $x$, the kk‑NN rule selects
@@
5.1 Excess‑risk of cost–spectrum k‑NN
@@
where $C$ depends only on $L$ and the diameter of $\Phi_{q}(\mathcal{X})$, and $r^{\star}$ is the Bayes‑optimal rule $r^{\star}(x)=\mathop{\rm arg\,min}_{m}\bigl[\gamma(x,h_{m})+\lambda c(h_{m})\bigr]$.
The proof follows the classical kk‑NN bound of [53, 5] with an extra $\lambda c(h)$ term that is constant w.r.t. $x$ and therefore preserves the rate.
@@ 5.2 Consistency of Cost–Spectrum InfoNCE
Fingerprints live on $\mathbb{S}^{d-1}$, so dot products are scaled similarities $S_{im}=\tfrac{q_{i}^{\top}e_{m}}{\tau_{k}}$ with band‑dependent temperature $\tau_{k}$ (Equation (7)). Let $\beta_{0}\!=\!0<\!\beta_{1}<\!\dots<\!\beta_{K}\!=\!1$ partition costs into bands $\mathcal{B}_{k}=\{m:\;c(h_{m})\!\in\![\beta_{k},\beta_{k+1})\}$, and denote $\mathcal{P}_{ik}=\mathcal{P}(i)\cap\mathcal{B}_{k}$ the correct experts for query $i$ that fall in band $k$. The Cost‑Spectrum InfoNCE loss for a single query $i$ is
@@ 5.3 Discussion
Theorem 5.2 guarantees that if query–expert descriptors are Lipschitz, CSCR converges to the Bayes‑optimal router at the usual kk‑NN rate. Lemma 5.3 justifies the specific form of our InfoNCE objective. Sec. 4 verifies these claims on three benchmarks. See proofs in Appendix B.
@@ 6 Conclusion
7 Acknowledgments
The authors would like to thank Zahra Miri for her assistance in preparing the figures.
This work was made possible by NSF IIS 2347592, 2348169, DBI 2405416, CCF 2348306, CNS 2347617.
@@ A.1 Enhancing and Optimizing LLMs
Large language models (LLMs) have demonstrated remarkable capabilities across diverse NLP tasks [66, 6]. To further improve their performance and efficiency, numerous strategies have been proposed.
@@
Enhancement approaches targeting individual LLMs include fine-tuning [67], prompting strategies like Chain-of-Thought (CoT) [102, 90], Tree-of-Thoughts [96], and inference-acceleration techniques such as early exiting [84, 103, 72] and speculative decoding [77, 83, 9, 45, 8]. Additionally, Mixture-of-Experts (MoE) architectures [36, 41, 75, 24, 104, 38] route inputs through sparse sub-models or "experts," reducing cost while retaining performance. However, these methods typically operate within a single LLM’s structure and may not generalize to multi-model scenarios.
@@
Fusion strategies synthesize outputs from multiple LLMs to improve output quality [68, 39, 29, 89, 52]. Fusion approaches often rely on unsupervised metrics [99, 73, 97] or ensemble voting to determine the final output [44]. A related but distinct technique is model merging [50], where weights from multiple pre-trained or fine-tuned models are combined—either directly via methods like weight averaging [92], Task Arithmetic [35], or Fisher merging [56]—or using more sophisticated techniques like TIES [94], AdaMerging [95], and ZipIt [78].
@@
In contrast, cascading invokes models sequentially—often ordered by computational cost—and halts once a satisfactory output is generated [10, 98, 10, 30]. Such approaches strike a balance between quality and efficiency, making them particularly attractive in production settings.
@@ A.2 LLM Routing
Non-predictive methods generate outputs from one or more models before making a selection. FrugalGPT [10] exemplifies this category, using a sequential strategy and a response quality threshold to minimize cost. Other works adopt layered inference architectures to escalate hard queries to more powerful models [91], or leverage cascades with self-verification [54, 98, 43, 51].
@@
In contrast, predictive routing aims to select the best model *before* any inference is performed. Strategies include supervised learning [76], reward-model-based routing [32, 51], and meta-models trained to predict LLM performance given an input [71]. Router models vary widely in implementation, including neural networks [19, 88, 11, 1], $k$-nearest neighbors [34, 76, 80, 43], matrix factorization [62, 105, 48], and graph neural networks [25]. Others incorporate model-specific tokens or train across multiple domains [18, 7].
@@
Routing and cascading are grounded in broader literature, including selective classification [27, 60], learning to defer [55], and learning to reject [12, 4, 14]. Several works explore supervision levels [51, 101], robustness [16, 59, 74], and evaluation frameworks for routers [33, 32].